Ammonia reacts with oxygen as shown below:

4 NH3 (g) + 5 O2 (g) → 4 NO (g) + 6 H2O (g) ΔH°rxn = -906 kJ

How much heat, in kJ, will be released by the production of 553 g of NO?

Give your answer in kJ, report your answer as a positive number, and do not use scientific notation.
1
Expert's answer
2020-04-28T12:21:57-0400

Solution:

The balanced chemical equation:

4NH3(g) + 5O2(g) → 4NO(g) + 6H2O(g), ΔH°rxn = -906 kJ 


Heat = Q = -ΔH°rxn = -(-906 kJ) = 906 kJ.


Moles of NO = Mass of NO / Molar mass of NO

Molar mass of NO is 30 g/mol. Then,

Moles of NO = n(NO) = (553 g) / (30 g/mol) = 18.433 mol

n(NO) = 18.433 mol


906 kJ of heat will be released by the production of 4 mol of NO (according to the chemical equation).

Q kJ of heat will be released by the production of 18.433 mol of NO (according to the task).

Therefore,

Q = (906 kJ) × (18.433 mol / 4 mol) = 4175.07 kJ = 4175 kJ


Answer: 4175 kJ of heat will be released by the production of 553 g of NO.